# Modeling individual attention dynamics on online social media

### v1

@@ Abstract
In an age characterized by an overwhelming abundance of information, the concept of the attention economy has emerged as a central framework for understanding how individuals navigate the digital landscape [1, 2, 3, 4]. Within this framework, attention, rather than information, is the truly scarce resource: Understanding how attention is captured, distributed, and ultimately decays, is crucial [5, 6, 7, 8]. Online social media platforms, for instance, are designed to facilitate interaction and information sharing, but bomb users with a constant stream of stimuli, including posts, comments, notifications, and messages [9]. Most of the previous research has focused on patterns of collective attention within these environments, such as the emergence of trending topics [10, 11, 12], information cascades [13], and the spread of memes [14, 15].
@@
However, individual-level attention dynamics, specifically how a single user’s focus evolves in response to multiple stimuli, remains mostly unexplored. This paper addresses this gap by introducing a simple model that describes the temporal decay of an individual’s attention when confronted with a multitude of interactive inputs. For example, consider a user who posts a thought-provoking question on a social media platform and receives numerous replies from different users, each demanding a piece of their attention [16, 17]. Similarly, the model could describe the attention dynamics in email inbox management: Individuals receive a stream of emails, each varying in importance, competing for limited attention [18, 19]. Likewise, in a collaborative work environment, a professional might receive numerous task assignments, comment threads, and status updates, and have to decide which assignment to prioritize [20, 21]. In all these examples, the model provides a framework to understand how an individual’s finite attention is distributed and how it decays when faced with a large number of inputs.
@@
To test our model, we focus on Reddit, a prominent online social news aggregation, content rating, and discussion platform. Reddit is organized into thousands of user-created communities known as subreddits, each dedicated to a specific topic. We specifically choose the subreddit Change My View (CMV) as our empirical testbed. Unlike many other subreddits, where users might post and then disengage, CMV is unique in its explicit goal: Users post an opinion and invite others to challenge it, actively encouraging and valuing engagement with the comments. Previous studies have focused on processing the text of comments in CMV, investigating the effects that the language structure can have on the power of persuasion in the discourse [22, 23, 24, 25] and on the process of reaching consensus [26]. Crucially, on CMV, users explicitly acknowledge comments that successfully “change their view”, allowing us to observe and quantify how individual attention evolves in response to various inputs. This inherent interest in following and responding to discussions makes CMV an ideal environment for studying individual attention decay, as users are demonstrably invested in the interactions surrounding their posts.
@@
Our model aims to predict the probability that an individual will allocate their attention to specific interactions over time. We derive an analytical solution for the attention decay as a function of the number of interactions received, their cumulative duration, and the distribution of activity among the interacting individuals. Interestingly, we found that the individual attention dynamics is fully determined by the overall duration of interactions, while independent of the number of interactions received and the users’ activity. Despite its simplicity, the model is able to capture the individual attention decay of users on CMV. Moreover, the model reproduces the interevent time between consecutive rewards, from both the point of view of the original post’s creator and comments’ authors. By modeling individual attention dynamics, we provide a microscopic perspective that complements the macroscopic view offered by collective attention studies [27, 28, 29, 30, 31].
@@
Comments in the thread are created by a population of $U$ users, with each user producing a total number of comments equal to $\eta$, obeying the distribution $P(\eta)$. If a comment draws the attention of the post’s creator, it can be rewarded with a positive acknowledgment. We refer to the authors of rewarded comments as receivers, while we refer to the probability that users are rewarded by their fitness $\rho$. We assume that the fitness $\rho$ is solely a function of the user’s activity $\eta$: The more comments a user posts, the more likely the user is rewarded. For simplicity, we consider $\rho\sim\eta^{\omega}$, with $\omega>0$. The probability that a randomly selected comment, written by a user with activity $\eta$, is rewarded is
@@
[…]
The goal of the model is to describe the attention dynamics of the post’s creator, which is manifested by comment rewarding. More specifically, we are interested in finding the days $t=t^{*}$ in which comments are rewarded, assuming that rewards are assigned the same day the comments are created. We characterize the probability distribution that a comment is rewarded on a certain day $t^{*}$, $P(t^{*})$. Such a probability is a function of the model’s parameters. We assume that the rate of reward $n_{0}$ is constant in time. Also, we assume that $N$ and $T$ are not correlated and are generated from the distributions $P(N)$ and $P(T)$, respectively. For each thread, we know the function $n(t)$ describing the number of comments made on day $t$; such a function is constrained by the actual size $N$ and duration $T$ of the thread. Finally, the population of users is fully described by the activity distribution $P(\eta)$ and the fitness exponent $\omega$. We first provide an analytical solution for $P(t^{*})$ where the above quantities are given by mathematical functions; next, we validate our model using empirical distributions obtained from the CMV dataset.
@@
For an exponential $n(t)$ given by Eq. (2), in the End Matter we show that $P(t^{*})$ is fully determined by the threads’ duration distribution $P(T)$ and is independent of the number of comments received, $P(N)$, and the activity of users, $P(\eta)$. As a particular case, we consider a power-law distribution $P(T)$ of the form
@@
with $T\in[T_{0},\infty)$, $B=(\delta-1)/T_{0}^{-(\delta-1)}$, and $\delta>2$. For this form of the $P(T)$, we obtain that, for $t^{*}\to 1$, the distribution $P(t^{*})$ goes to a constant if $\lambda\ll T_{0}$, while for $t^{*}\to\infty$ we get the asymptotic behavior $P(t^{*})\sim(t^{*})^{-\delta}$ (see End Matter for details). The main panel of Fig. 1(a) shows the distribution $P(t^{*})$ obtained from numerical simulations, which confirms the predicted behaviors. Importantly, to get a constant $P(t^{*})$ for low values of $t^{*}$, we ensured that the condition $\lambda\ll T_{0}$ is satisfied.
Data validation. We empirically validate our model by measuring how individuals allocate their attention on CMV, where users reward their peers for interesting comments. On CMV, users submit posts expressing a certain opinion, and others respond to the post, attempting to change the poster’s view by providing counterarguments. The original post’s creator acknowledges the relevant comments by rewarding the author of the successful comment. We use these rewards as traces to monitor the attention dynamics of the post’s creator, i.e., to reconstruct the probability distribution $P(t^{*})$ that a comment is rewarded at time $t^{*}$.
We collected data from CMV comprising $V=102296$ threads and $C=6564271$ comments posted by $U=427742$ users from its creation in 2013 to 2023, see End Matter for details. Each thread $i$ is characterized by the number of comments $n_{i}(t)$ posted on day $t$ since the creation of the thread. The duration of the thread $i$ is denoted by $T_{i}$, which means that $n_{i}(t)=0$ for $t>T_{i}$; the total size of the thread $i$ is then $N_{i}=\sum_{t}n_{i}(t)$. We empirically observe that $1\leq T_{i}\leq 180$ for all threads, indicating that comments are not allowed more than 6 months after the post’s creation date. We define the empirical distributions of thread duration as $P(T)=\frac{1}{V}\sum_{i=1}^{V}\delta_{T,T_{i}}$ and thread size as $P(N)=\frac{1}{V}\sum_{i=1}^{V}\delta_{N,N_{i}}$, where the Kronecker delta function is such that $\delta_{x,y}=1$ if $x=y$ and $\delta_{x,y}=0$ otherwise. Further, each user $u$ is characterized by the total number of comments posted $\eta_{u}$ and the total number of rewarded comments $0\leq\eta^{*}_{u}\leq\eta_{u}$. The empirical distribution of users’ activity is defined as $P(\eta)=\frac{1}{U}\sum_{u=1}^{U}\delta_{\eta,\eta_{u}}$.
@@
For the $i$-th thread, the conditional probability of observing a reward on day $t=t^{*}$, given $N_{i}$ and $T_{i}$, is defined as (see End Matter)
@@
Inter-reward dynamics. After investigating the probability that a comment is rewarded on a certain day, $P(t^{*})$, we test if the model is able to capture the inter-reward dynamics. Even though $P(t^{*})$ is independent of the users’ fitness $\rho$, the fitness distribution $P(\rho)$ is central to determining the probability that a user is rewarded. We can measure the fitness of a user as the number of times that has been rewarded in total, in such a way that the fitness of receivers is nonzero. In Fig. SF1 of the Supplementary Material (SM) we show the number of rewarded comments, $\eta^{*}$, as a function of the number of comments posted, $\eta$, for every receiver in CMV. As one could expect, there is a clear correlation between both quantities: The more (less) comments a user posts the more (less) rewards receives. However, since the dispersion is non-negligible, we cannot assume an analytical relationship of the form $\rho\sim\eta^{\omega}$ by directly performing a nonlinear regression. Instead, we can sample $\rho$ from the posterior distribution $P(\rho|\eta)$ obtained via the Bayes’ rule [33]. First, we approximate the conditional probability $P(\eta|\rho)$ of observing $\eta$ given $\rho$ as normal distribution with mean $\mu$ and standard deviation $\sigma$ given by
@@
with $\nu_{\mu}=0.91$ and $\nu_{\sigma}=0.68$, see regression analysis in Fig. SF2 of the SM. Considering that the prior probability $P(\rho)$ is defined by the distribution of rewarded comments posted by a single user in CMV, $P(\eta^{*})\sim(\eta^{*})^{-\gamma_{\rho}}$, we can write the posterior distribution as
@@
We feed our mathematical model using spline interpolations [34] of the empirical distributions $P(N)$, $P(T)$, and $P(\eta)$, see Fig. SF3 of the SM. In Fig. 2(a) we compare the distribution of rewarded comments $P(\eta^{*})$ obtained in the model with the one directly measured in the data. We observe that the model reproduces the real distribution with good accuracy, showing a power-law signature $P(\eta^{*})\sim(\eta^{*})^{-\gamma_{\rho}}$ with exponents $\gamma_{\rho}=2.19\pm 0.02$ for the data and $\gamma_{\rho}=1.86\pm 0.07$ for the model. It should be stressed that, according to the empirical data, we restrict the possible number of rewards per user in the same thread to 1. This means that if a receiver has 100 rewards in CMV, they must be rewarded in 100 different threads.
@@
Figure 2: (a) Probability distribution $P(\eta^{*})$ of number of comments posted by a single user that have been rewarded. (b) Probability distribution $P(\tau)$ of elapsed days between two consecutive rewarded comments in threads. (c) Main: Probability distribution $P(m)$ of elapsed comments between two consecutive rewards of a single user. Insets: Nonlinear regression of a stretched exponential performed in $P(m)$ with exponents $\beta=0.52\pm 0.02$ (CMV) and $\beta=0.56\pm 0.02$ (model). We show the results obtained from both CMV and the model, represented by red and blue symbols, respectively. We used $n_{0}=200$ in Eq. (1).
@@
We also study the inter-reward dynamics by counting the number of days between two consecutive rewarded comments. By definition, the first possible reward in a thread is given in $t=1$. Therefore, if, for example, the first rewarded comment is posted on day $t=1$, the number of elapsed days is $\tau=0$. If the second rewarded comment is posted on $t=5$, then $\tau=4$. We compare the distribution of elapsed days $P(\tau)$ observed empirically in CMV with the one obtained by numerical simulations of the model. To this aim, in the model we consider $V=10^{5}$ threads and $U=4.5\times 10^{5}$ users. Figure 2(b) shows the interevent time distribution $P(\tau)$ in CMV (red symbols) and in the model (blue symbols). We observe that the model reproduces the real distribution with good accuracy, showing a power-law signature $P(\tau)\sim\tau^{-\gamma_{t}}$ with exponents $\gamma_{t}=2.30\pm 0.08$ for the data and $\gamma_{t}=2.02\pm 0.17$ for the model.
@@
Finally, we study the inter-reward dynamics from the point of view of the comments’ authors. For each receiver, i.e., a user who has been rewarded at least once, we define the time-ordered series of comments they authored on different threads. Then, we express the elapsed time between two consecutive rewards as the number of comments occurred between them. Again, the first possible reward is given in the first position in the series. Therefore, if, for example, the first reward is given to the first comment, the number of elapsed comments is $m=0$. If the second reward is given to the fifth comment, then $m=4$. The main panel of Fig. 2(c) shows the distribution of elapsed comments $P(m)$ observed empirically in CMV (red symbols) and obtained numerically from the model (blue symbols). We observe that the model accurately reproduces the real distribution. In the insets of Fig. 2(c), we show that the distribution $P(m)$ follows a stretched exponential $P(m)\sim\exp{\left(-a\,m^{\beta}\right)}$. By performing a nonlinear regression (solid lines), we obtain stretching exponents $\beta=0.52\pm 0.02$ for the data and $\beta=0.56\pm 0.02$ for the model.
@@
Our main contribution is a minimal, analytically tractable model that captures how individual attention decays over time, as a function of the total duration of interactions, rather than their number or the activity levels of interacting users. We validated the model using empirical data from the subreddit Change My View, where posts’ creators engage with comments by rewarding them—an observable proxy for individual attention. Despite its simplicity, the model successfully reproduced both the decay in attention of the original post’s creator and the interevent times between successive rewards. These findings underscore the model’s predictive power and its potential applicability to a wide range of online contexts, from email management to collaborative work platforms.
The model’s simplicity trades off realism for analytical tractability, which opens several avenues for future research. For instance, one could incorporate heterogeneity in user fitness, linking it not only to activity but also to an intrinsic capacity to steer focal user’s attention. In this setting, the individual attention dynamics is likely to depend also on the heterogeneity of the fitness distribution. Furthermore, one could validate our model on other data sets describing distinct social contexts, where the individual attention dynamics may be driven by different mechanisms.

### v2

@@ Abstract
In an age characterized by an overwhelming abundance of information, the concept of the attention economy has emerged as a central framework for understanding how individuals navigate the digital landscape [24, 25, 7, 31]. Within this framework, attention, rather than information, is the truly scarce resource: Understanding how attention is captured, distributed, and ultimately decays, is crucial [48, 23, 11, 36, 28]. Online social media platforms, for instance, are designed to facilitate interaction and information sharing, but bomb users with a constant stream of stimuli, including posts, comments, notifications, and messages [37]. Most of the previous research has focused on patterns of collective attention within these environments, such as the emergence of trending topics [3, 2, 21], information cascades [40, 19], and the spread of memes [46, 20].
@@
However, individual-level attention dynamics, specifically how a single user’s focus evolves in response to multiple stimuli, remains mostly unexplored. This paper addresses this gap by introducing a simple model that describes the temporal decay of an individual’s attention when confronted with a multitude of interactive inputs. For example, consider a user who posts a thought-provoking question on a social media platform and receives numerous replies from different users, each demanding a piece of their attention [42, 16]. Similarly, the model could describe the attention dynamics in email inbox management: Individuals receive a stream of emails, each varying in importance, competing for limited attention [47, 5, 45]. Likewise, in a collaborative work environment, a professional might receive numerous task assignments, comment threads, and status updates, and have to decide which assignment to prioritize [30, 8]. In all these examples, the model provides a framework to understand how an individual’s finite attention is distributed and how it decays when faced with a large number of inputs.
@@
To test our model, we focus on Reddit, a prominent online social news aggregation, content rating, and discussion platform. Reddit is organized into thousands of user-created communities known as subreddits, each dedicated to a specific topic. We specifically choose the subreddit Change My View (CMV) as our empirical testbed. Unlike many other subreddits, where users might post and then disengage, CMV is unique in its explicit goal: Users post an opinion and invite others to challenge it, actively encouraging and valuing engagement with the comments. Previous studies have focused on processing the text of comments in CMV, investigating the effects that the language structure can have on the power of persuasion in the discourse [41, 35, 32, 14] and on the process of reaching consensus [34]. Furthermore, CMV has been widely used to study users’ opinion change on social media [41, 32, 34], among other online platforms [49]. Crucially, on CMV, users explicitly acknowledge comments that successfully “change their view”, allowing us to observe and quantify how individual attention evolves in response to various inputs. This inherent interest in following and responding to discussions makes CMV an ideal environment for studying individual attention decay, as users are demonstrably invested in the interactions surrounding their posts.
@@
Our model aims to predict the probability that an individual will allocate their attention to specific interactions from others, over time. The model assumes that the focal individual rewards their peers with attention simply based on their activity in writing comments. We derive an analytical solution for the attention decay as a function of the number of interactions received, their cumulative duration, and the distribution of activity among the interacting individuals. We found that the individual attention dynamics is fully determined by the overall duration of interactions, while independent of the number of interactions received and the users’ activity. Despite its simplicity, the model is able to capture the individual attention decay of users on CMV. Moreover, the model reproduces the interevent time between consecutive rewards, from both the point of view of the original poster (OP) and comments’ authors (or commenters). By modeling individual attention dynamics, we provide a microscopic perspective that complements the macroscopic view offered by collective attention studies [27, 38, 15, 33, 4, 17, 22, 29, 13, 6, 26].
@@
Comments in the thread are created by a population of $U$ users, with each user producing a total number of comments equal to $\eta$, obeying the distribution $P(\eta)$. If a comment draws the attention of the OP, it can be rewarded with a positive acknowledgment. We refer to the authors of rewarded comments as receivers, while we refer to the probability that users are rewarded by their fitness $\rho$. We assume that the fitness $\rho$ is solely a function of the user’s activity $\eta$: The more comments a user posts, the more likely the user is rewarded. For simplicity, we consider $\rho\sim\eta^{\omega}$, with $\omega>0$. The probability that a randomly selected comment, written by a user with activity $\eta$, is rewarded is
@@
where $n_{0}$ is the rate of rewarded comments and $\langle\ldots\rangle$ denotes the mean value, see Appendix for details. Importantly, Eq. (1) is valid for a single comment. In case of $n(t)$ multiple comments posted on day $t$, the probability that at least one of the comments written by a user with activity $\eta$ is rewarded is given by $1-\left[1-p_{r}(\eta)\right]^{n(t)}$.
The goal of the model is to describe the attention dynamics of the OP, which is manifested by comment rewarding. More specifically, we are interested in finding the days $t=t^{*}$ in which comments are rewarded, assuming that rewards are assigned the same day the comments are created. We characterize the probability distribution that a comment is rewarded on a certain day $t^{*}$, $P(t^{*})$. Such a probability is a function of the model’s parameters. We assume that the rate of reward $n_{0}$ is constant in time. Also, we assume that $N$ and $T$ are not correlated and are generated from the distributions $P(N)$ and $P(T)$, respectively. For each thread, we know the function $n(t)$ describing the number of comments made on day $t$; such a function is constrained by the actual size $N$ and duration $T$ of the thread. Finally, the population of users is fully described by the activity distribution $P(\eta)$ and the fitness exponent $\omega$. We first provide an analytical solution for $P(t^{*})$ where the above quantities are given by mathematical functions; next, we validate our model using empirical distributions obtained from the CMV dataset.
@@
For an exponential $n(t)$ given by Eq. (2), in Appendix we show that $P(t^{*})$ is fully determined by the threads’ duration distribution $P(T)$ and is independent of the number of comments received, $P(N)$, and the activity of users, $P(\eta)$. As a particular case, we consider a power-law distribution $P(T)$ of the form
@@
with $T\in[T_{0},\infty)$, $B=(\delta-1)/T_{0}^{-(\delta-1)}$, and $\delta>2$. For this form of the $P(T)$, we obtain that, for $t^{*}\to 1$, the distribution $P(t^{*})$ goes to a constant if $\lambda\ll T_{0}$, while for $t^{*}\to\infty$ we get the asymptotic behavior $P(t^{*})\sim(t^{*})^{-\delta}$ (see Appendix for details). The main panel of Fig. 1(a) shows the distribution $P(t^{*})$ obtained from numerical simulations, which confirms the predicted behaviors. Importantly, to get a constant $P(t^{*})$ for low values of $t^{*}$, we ensured that the condition $\lambda\ll T_{0}$ is satisfied. We observe that the crossover between the two predicted regimes—constant for $t^{*}\to 1$ and power-law for $t^{*}\to\infty$—occurs around $t^{*}=T_{0}$.
Data validation. We empirically validate our model by measuring how individuals allocate their attention on CMV, where users reward their peers for interesting comments. On CMV, users submit posts expressing a certain opinion, and others respond to the post, attempting to change the OP’s view by providing counterarguments. The OP acknowledges the relevant comments by rewarding the author of the successful comment. We use these rewards as traces to monitor the attention dynamics of the OP, i.e., to reconstruct the probability distribution $P(t^{*})$ that a comment is rewarded at time $t^{*}$.
We collected data from CMV comprising $V=102296$ threads and $C=6564271$ comments posted by $U=427742$ users from its creation in 2013 to 2023, see Section I of the Supplementary Material for details. Each thread $i$ is characterized by the number of comments $n_{i}(t)$ posted on day $t$ since the creation of the thread. The duration of the thread $i$ is denoted by $T_{i}$, which means that $n_{i}(t)=0$ for $t>T_{i}$; the total size of the thread $i$ is then $N_{i}=\sum_{t}n_{i}(t)$. We empirically observe that $1\leq T_{i}\leq 180$ for all threads, indicating that comments are not allowed more than 6 months after the post’s creation date. We define the empirical distributions of thread duration as $P(T)=\frac{1}{V}\sum_{i=1}^{V}\delta_{T,T_{i}}$ and thread size as $P(N)=\frac{1}{V}\sum_{i=1}^{V}\delta_{N,N_{i}}$, where the Kronecker delta function is such that $\delta_{x,y}=1$ if $x=y$ and $\delta_{x,y}=0$ otherwise. Further, each user $u$ is characterized by the total number of comments posted $\eta_{u}$ and the total number of rewarded comments $0\leq\eta^{*}_{u}\leq\eta_{u}$. The empirical distribution of users’ activity is defined as $P(\eta)=\frac{1}{U}\sum_{u=1}^{U}\delta_{\eta,\eta_{u}}$.
@@
For the $i$-th thread, the conditional probability of observing a reward on day $t=t^{*}$, given $N_{i}$ and $T_{i}$, is defined as (see Appendix)
@@
Inter-reward dynamics. After investigating the probability that a comment is rewarded on a certain day, $P(t^{*})$, we test if the model is able to capture the inter-reward dynamics. Even though $P(t^{*})$ is independent of the users’ fitness $\rho$, the fitness distribution $P(\rho)$ is central to determining the probability that a user is rewarded. We can measure the fitness of a user as the number of times that has been rewarded in total, in such a way that the fitness of receivers is nonzero. In Supplementary Fig. 1 we show the number of rewarded comments, $\eta^{*}$, as a function of the number of comments posted, $\eta$, for every receiver in CMV. As one could expect, there is a clear correlation between both quantities: The more (less) comments a user posts the more (less) rewards receives. However, since the dispersion is non-negligible, we cannot assume an analytical relationship of the form $\rho\sim\eta^{\omega}$ by directly performing a nonlinear regression. Instead, we can sample $\rho$ from the posterior distribution $P(\rho|\eta)$ obtained via the Bayes’ rule [18]. First, we approximate the conditional probability $P(\eta|\rho)$ of observing $\eta$ given $\rho$ as normal distribution with mean $\mu$ and standard deviation $\sigma$ given by
@@
with $\nu_{\mu}=0.91$ and $\nu_{\sigma}=0.68$, see regression analysis in Supplementary Fig. 2. Considering that the prior probability $P(\rho)$ is defined by the distribution of rewarded comments posted by a single user in CMV, $P(\eta^{*})\sim(\eta^{*})^{-\gamma_{\rho}}$, we can write the posterior distribution as
@@
We feed our mathematical model using spline interpolations [44] of the empirical distributions $P(N)$, $P(T)$, and $P(\eta)$, see Supplementary Fig. 3. Similarly, we use the thread-specific time series given by Eq. (4) as the number of comments throughout the days. In Fig. 2(a) we compare the distribution of rewarded comments $P(\eta^{*})$ obtained in the model with the one directly measured in the data. We observe that the model reproduces the real distribution with good accuracy, showing a power-law signature $P(\eta^{*})\sim(\eta^{*})^{-\gamma_{\rho}}$ with exponents $\gamma_{\rho}=2.19\pm 0.02$ in the data and $\gamma_{\rho}=1.86\pm 0.07$ in the model. It should be stressed that, according to the empirical data, we restrict the possible number of rewards per user in the same thread to 1. This means that if a receiver has 100 rewards in CMV, they must be rewarded in 100 different threads.
@@
Figure 2: (a) Probability distribution $P(\eta^{*})$ of number of comments posted by a single user that have been rewarded. (b) Probability distribution $P(\tau)$ of elapsed days between two consecutive rewarded comments in threads. (c) Main: Probability distribution $P(m)$ of elapsed comments between two consecutive rewards of a single user. Insets: Nonlinear regression of a stretched exponential performed in $P(m)$ with exponents $\beta=0.52\pm 0.02$ (CMV) and $\beta=0.56\pm 0.02$ (model). We show the results obtained from both CMV and the model, represented by red and blue symbols, respectively. Very similar results are obtained by running the model with different initial conditions. We used a logarithmic binning where bins are sized like the Fibonacci sequence [43]. We used $n_{0}=200$ in Eq. (1).
@@
We also study the inter-reward dynamics by counting the number of days between two consecutive rewarded comments. By definition, the first possible reward in a thread is given in $t=1$. Therefore, if, for example, the first rewarded comment is posted on day $t=1$, the number of elapsed days is $\tau=0$. If the second rewarded comment is posted on $t=5$, then $\tau=4$. We compare the distribution of elapsed days $P(\tau)$ observed empirically in CMV with the one obtained by numerical simulations of the model. To this aim, in the model we consider $V=10^{5}$ threads and $U=4.5\times 10^{5}$ users. Figure 2(b) shows the interevent time distribution $P(\tau)$ in CMV (red symbols) and in the model (blue symbols). We observe that the model reproduces the real distribution with good accuracy, showing a power-law signature $P(\tau)\sim\tau^{-\gamma_{t}}$ with exponents $\gamma_{t}=2.30\pm 0.08$ in the data and $\gamma_{t}=2.02\pm 0.17$ in the model.
@@
Finally, we study the inter-reward dynamics from the point of view of the commenters. For each receiver, i.e., a user who has been rewarded at least once, we define the time-ordered series of comments they authored on different threads. Then, we express the elapsed time between two consecutive rewards as the number of comments occurred between them. Again, the first possible reward is given in the first position in the series. Therefore, if, for example, the first reward is given to the first comment, the number of elapsed comments is $m=0$. If the second reward is given to the fifth comment, then $m=4$. The main panel of Fig. 2(c) shows the distribution of elapsed comments $P(m)$ observed empirically in CMV (red symbols) and obtained numerically from the model (blue symbols). We observe that the model accurately reproduces the real distribution. In the insets of Fig. 2(c), we show that the distribution $P(m)$ follows a stretched exponential $P(m)\sim\exp{\left(-a\,m^{\beta}\right)}$. By performing a nonlinear regression (solid lines), we obtain stretching exponents $\beta=0.52\pm 0.02$ in the data and $\beta=0.56\pm 0.02$ in the model.
@@
Our main contribution is a minimal, analytically tractable model that captures how individual attention decays over time, as a function of the total duration of interactions, rather than their number or the activity levels of interacting users. We validated the model using empirical data from the subreddit Change My View, where OPs engage with comments by rewarding them—an observable proxy for individual attention. Despite its simplicity, the model successfully reproduced both the decay in OP’s attention and the interevent times between successive rewards. These findings underscore the model’s predictive power and its potential applicability to a wide range of online contexts, from email management to collaborative work platforms.
The model’s simplicity trades off realism for analytical tractability, which opens several avenues for future research. From a theoretical point of view, the model proposed can be interpreted as a stochastic counting process of attention allocation, susceptible to an application of the tools of renewal theory [12] to obtain further information. From the methodological side, on the other hand, heterogeneity could be incorporated in user fitness, linking it not only to activity but also to an intrinsic capacity to steer focal user’s attention, content-based preference, or interaction similarity between users. In this setting, the individual attention dynamics is likely to depend also on the heterogeneity of the fitness distribution. Furthermore, one could validate our results on other data sets describing distinct social contexts, where the individual attention dynamics may be driven by different mechanisms. The generality of the model can be tested by considering alternative online platforms different from Reddit. Finally, it has been recently shown that individual attention is closely related to how users express their emotions on social media [50]. Similarly to our work, future efforts should extend the framework of collective emotions to a microscopic, individual level [39, 10].
Acknowledgments. J. O., S. F., and M. S. acknowledge the support of the AccelNet-MultiNet program, a project of the National Science Foundation (award #1927425 and #1927418). J. O. and R. P.-S. acknowledge financial support from project PID2022-137505NB-C21/AEI/10.13039/501100011033/FEDER UE. F. R. acknowledges support from the Air Force Office of Scientific Research under grant number FA9550-24-1-0039. M. S. acknowledges support from Grants No. RYC2022-037932-I and CNS2023-144156 funded by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. We thank the Indiana University Observatory on Social Media (OSoMe) for managing and providing access to Reddit data. We also acknowledge Jisun An and Haewoon Kwak for initially bringing this dataset to OSoMe’s attention.